OpenStudy (anonymous):

Use a graphing calculator to solve the equation 3tan 1/3 0=8 in the interval from 0 to 2pi Round your answers to the nearest hundredth.

jimthompson5910 (jim_thompson5910):

This equation? \[\Large 3\tan\left(\frac{1}{3}\theta\right) = 8\]

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

ok replace theta with x \[\Large 3\tan\left(\frac{1}{3}\theta\right) = 8\] \[\Large 3\tan\left(\frac{1}{3}x\right) = 8\]

jimthompson5910 (jim_thompson5910):

then subtract 8 from both sides \[\Large 3\tan\left(\frac{1}{3}x\right) = 8\] \[\Large 3\tan\left(\frac{1}{3}x\right)-8 = 8-8\] \[\Large 3\tan\left(\frac{1}{3}x\right)-8 = 0\]

jimthompson5910 (jim_thompson5910):

now graph `y = 3*tan(1/3x)-8` into a graphing calculator here's a free one to use if you don't have a graphing calculator https://www.desmos.com/calculator

jimthompson5910 (jim_thompson5910):

look for the point where the graph crosses the x axis

OpenStudy (anonymous):

this is what i am given

jimthompson5910 (jim_thompson5910):

what did you type in?

OpenStudy (anonymous):

y=3*tan (1/3x)-8

jimthompson5910 (jim_thompson5910):

strange, I'm getting this (see attached image)

OpenStudy (campbell_st):

if you are being asked to solve, why not graph y = 8 and \[y = \tan(\frac{x}{3})\] and find the points of intersection in the given domain.

jimthompson5910 (jim_thompson5910):

@campbell_st it should be `3tan(x/3)`

jimthompson5910 (jim_thompson5910):

but that's a good alternative

OpenStudy (anonymous):

that is weird mine will not give me that graph. so approx -5.9 and 3.7

jimthompson5910 (jim_thompson5910):

https://www.desmos.com/calculator/fgkktgo5bo

jimthompson5910 (jim_thompson5910):

focus on the interval from 0 to 2pi 2pi = 6.28 so you're only focusing on roots between 0 and 6.28

OpenStudy (anonymous):

3.64 is the only point located between the two so that would be the answer

jimthompson5910 (jim_thompson5910):

3.64 is correct

OpenStudy (anonymous):

Thank you again for explaining!

jimthompson5910 (jim_thompson5910):

sure thing

OpenStudy (campbell_st):